# SkillDiffuser: Interpretable Hierarchical Planning via Skill Abstractions in Diffusion-Based Task Execution

### v1

@@ 4.4 Training the SkillDiffuser
We engineered a threefold loss function, with each meticulously designed and straightforward to meet the distinct needs of respective modules within SkillDiffuser.
For the task-specific inverse dynamics model, we utilize a behavior cloning loss [42] to train our inverse MLP which emulates expert actions from observed state transitions. This loss, denoted as $\mathcal{L}_{inv}$, is formulated as:
$\mathcal{L}_{inv}=\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\left[\left\|\boldsymbol{a}-\Psi(\tilde{\boldsymbol{s}},\tilde{\boldsymbol{s}}^{\prime},\boldsymbol{i})\right\|_{2}^{2}\right],$ (10)
@@
Correspondingly, the other parts of the proposed model are all task-agnostic. For high-level skill abstraction module, we consider the vector quantization (VQ) loss to optimize the skill predictor. The VQ loss, $\mathcal{L}_{\mathrm{VQ}}$, encourages the skill predictor embeddings to adhere to the skill set vectors, enhancing the interpretability and consistency of the skill representations. Inspired by VQ-VAE [43], we have
$\mathcal{L}_{\mathrm{VQ}}=\mathbb{E}_{\boldsymbol{\tau}}\left[\|\mathbf{q}(\tilde{\boldsymbol{z}})-\tilde{\boldsymbol{z}}\|_{2}^{2}\right],$ (11)
@@
Table 1: Task-wise success rates on LOReL Sawyer Dataset. We show our success rates compared to random policy, language-conditioned imitation learning (LCBC), language-conditioned Q-learning (LCRL), a flat non-hierarchical Decision Transformer (Lang-DT), and LISA. The results on each dataset are calculated over 3 seeds. SkillDiffuser outperforms all other methods in terms of average performance over all tasks. Best methods and those within 5% of the best are highlighted in bold.
@@
Additionally, the language encoder is achieved by leveraging a pre-trained GPT-2 [34] with parameters frozen to ensure stability in the language understanding component. And we employ diverse settings to serve as the visual encoder to ensure fair comparison in different datasets. We elaborate the details in corresponding parts of Sec. 5. More training details are shown in Appendix E and we also provide some pseudo-code of our algorithm in Appendix B.
@@ 5.1 Datasets
LOReL Sawyer Dataset [28]. The Language-conditioned Offline Reward learning dataset, abbreviated as LOReL, is composed of pseudo-expert trajectories or play data gathered from the replay buffer of an arbitrary reinforcement learning policy, annotated with post-hoc crowd-sourced language directives. The inherent randomness within these trajectories presents a significant challenge for behavior cloning applications. The LOReL Sawyer dataset encompasses 50k trajectories, each with 20 steps, in a simulated Sawyer robot environment. We assess our approach using the same six tasks as the original paper [28]. This comes to a total of 77 instructions for all 6 tasks combined. More details about the dataset can be found in Appendix D.1.
@@ 5.2 Evaluation Results on LOReL Sawyer Dataset
Settings. We first perform experiments on LOReL Sawyer dataset with only partially observed image data. The analysis is anchored on six different tasks with paraphrased instructions under five different situations (i.e. unseen verb, unseen none, human provided, etc. as introduced in Sec. 5.1). This deliberate choice facilitates a direct comparison of our method against established benchmarks.
Baselines. For our baselines, we employ random policy, language-conditioned imitation learning (LCBC) [40], language-conditioned Q-learning (LCRL) [19], Lang-DT (which is called Flat Baseline in [12]) and state-of-the-art skill-learning based method LISA [12]. We follow the settings of LOReL [28] paper for the first three algorithms, while follow LISA [12] for the last two.
The random policy serves as a baseline, executing actions without specific strategy or learning and providing a fundamental comparison point for the algorithm effectiveness. LCBC mimics offline dataset behaviors based on language instructions, aligning with previous works focusing on imitation learning to achieve language-conditioned behavior. In contrast, LCRL employs reinforcement learning, labeling each episode’s final state with language-instructed rewards. And Lang-DT serves as a non-hierarchical benchmark that utilizes a Causal Transformer [24] with language instructions. Moreover, LISA incorporates a dual-transformer structure, with one for skill prediction and the other for action planning, achieving state-of-the-art performance in language instructed multi-task learning.
@@
Method Lang Diffuser Skill Set Performance
@@
Table 3: Ablation study of language skill conditioning on Meta-World. Results of Flat R3M, Language-condition diffuser and our SkillDiffuser are shown here. All are averaged over 3 runs.
5.3 Ablation Study on Meta-World Dataset
Settings. We conduct experiments on Meta-World MT10 benchmark, a standard multi-task dataset which contains ten carefully selected tasks with sub-optimal expert data. We still do not use state data of the robotics but exploit visual observations only. We finely annotate the ten tasks with language instructions. Details are shown in Appendix D.2.
Baselines. We evaluate our methods against two baselines, both modifications of existing models. The first, Flat R3M, is adapted from the R3M paper [29]’s planner. As the original one utilizes the first four terms of robotic state information (i.e. proprio = 4 in R3M implementation), we reset “proprio” to zero, focusing exclusively on visual observations, and train the model on a multi-task dataset with 100 demonstrations per task (1000 total). To ensure fairness in comparison, we use R3M as our visual encoder in this setting. The second baseline is a variant of our SkillDiffuser, lacking the high-level skill abstraction module but retaining the low-level conditional diffusion-based planner, to assess the impact of skill abstraction. This version integrates a two-layer MLP to predict options from visual and language inputs, functioning as a language-conditioned diffuser, thus preserving the original setting’s core elements.
@@ 5.3 Ablation Study on Meta-World Dataset
Ablation on Language Skill Conditioning. Table 3 indicates that the Meta-World MT10 tasks, reliant solely on visual inputs without state information, are challenging. The Flat R3M method [29], lacking language conditioning and skill sets, achieves moderate performance on tasks like drawer-close and reach just through behavior cloning. Traditional Lang Diffuser, with added language conditioning, outperforms Flat R3M. This suggests the value of language elements in enhancing performance. Unlike these diffusers that treat the entire task space as a single skill, our SkillDiffuser, by discretizing skills into a skill set, achieves a 6% higher performance than language-conditioned diffusers, demonstrating the effectiveness of skill acquisition.
5.4 Visualization Results of Skill Set
@@
We show the visualization results of skill set on Meta-World Multi-Task 10 (MT10) here in Fig. 4 and result on LOReL dataset in Fig. 5 of Appendix C. The visual analysis of our SkillDiffuser’s learned skills on MT10 benchmark within Meta-World reveals that out of the 20-size skill set, our method learned eight skills (e.g. open door with revolving joint, pick and place goal etc.) notably distinguished by their unique word highlights. These bright spots across eight columns in the heatmap underscore the model’s ability to isolate and identify distinct skills from visual input, without an explicitly defined skill library. This indicates not only a significant interpretative advancement over traditional diffusion model-based planning but also the successful abstraction of high-level skill representations that are comprehensible to humans.
@@
Figure 4: Visualization result of skill heat map on Meta-World Multi-Task 10 (MT10). We display the word frequency associated with a skill set of size 20 in Meta-World MT10, normalized by column. The data’s sparsity and distinct highlights indicate certain language tokens are uniquely linked to specific skills. There are eight skills learned by our method. (zoom in for best view)
@@ 6 Conclusion
This paper presents SkillDiffuser, an integrated framework that enables robots to execute tasks from natural language instructions by enabling interpretable skill learning and conditional diffusion planning. It uses vector quantization to learn discrete yet understandable skill representations directly from visual and linguistic demonstrations. These skills condition a diffusion model to generate state trajectories adhering to the learned skills. By combining hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can follow abstract instructions to complete manipulation tasks. Extensive Experiments on manipulation benchmarks demonstrate state-of-the-art performance, highlighting its effectiveness for long-horizon tasks and ability to automatically learn interpretable skills.
@@ Appendix C More Visualizations
C.1 Visualization Results of Skill Set
Heat Map of Word Frequency in LOReL Dataset. As mentioned before, we show the visualization results of skill set on LOReL Sawyer Dataset in Fig. 5. The visualization results show that out of a 20-size skill-set, our SkillDiffuser learned 11 skills (e.g. shut close container drawer, pull drawer handle etc.) notably distinguished by their unique word highlights. The results demonstrate strong skill abstraction abilities. For example, the skill “shut close container drawer” abstracts different expressions like “shut drawer”, “shut container” into one skill semantic. In the heatmap, the presence of distinct bright spots across eleven columns strongly reaffirms the model’s capability to discern and pinpoint specific skills from visual inputs, in the absence of a pre-defined skill library. This observation is not just a testament to the model’s enhanced interpretative prowess over conventional diffusion-based planning approaches but also marks a remarkable stride in abstracting high-level skills into representations that are intuitively understandable by humans. Such evidence further validates the model’s proficiency in sophisticated skill identification and representation.
@@ C.2 Word Cloud of Learned Skills
Task Description
Closing the Drawer
Involves the robot’s precise manipulation of a drawer to close it, testing spatial dynamics understanding and fine motor control.
Turning the Faucet Left
Assesses the robot’s precision in rotational movements for turning a faucet to the left, a nuanced everyday action.
Pushing the Black Mug Right
Requires the robot to push a specific object (black mug) to the right, testing its skills in object recognition and directional movement.
Table 4: Overview of tasks in LOReL Sawyer Dataset.
@@ D.1 LOReL Sawyer Dataset
D.2 Meta-World Dataset
We sample 100 trajectories for each task of MT10 and form the expert dataset of 1000 trajectories.
@@ D.2 Meta-World Dataset
Table 5: Our annotated language instructions of Meta-World MT10 tasks.
Appendix E Implementation Details
E.1 Hyper-parameters
@@ E.2 Architecture Details
Figure 9: Task-wise success rates (in %) on LOReL Sawyer Dataset.
E.3 Training Details
@@
Figure 12: Task-wise success rates (in %) on Meta-World MT10 Dataset.
Appendix F More Results
F.1 Task-wise Performance on LOReL Dataset
We further demonstrate the performance of our method and other baselines on LOReL Sawyer dataset in Fig 9 and 10. As can be seen from the figures, especially from Fig. 10, our method’s average performance on 5 rephrases is nearly 10 percentage points higher than the previous SOTA, which demonstrates its strong robustness against ambiguous language instructions.
F.2 Task-wise Performance on Meta-World Dataset

### v3

@@ 4.4 Training the SkillDiffuser
We engineered a threefold loss function for SkillDiffuser. Firstly, for the inverse dynamics model which is task-specific, we employ a behavior cloning loss [44] to train our inverse MLP emulating expert actions from observed state transitions. This loss, denoted as $\mathcal{L}_{inv}$, is formulated as:
$\mathcal{L}_{inv}=\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D}}\left[\left\|\boldsymbol{a}-\Psi(\tilde{\boldsymbol{s}},\tilde{\boldsymbol{s}}^{\prime},\boldsymbol{i})\right\|_{2}^{2}\right],\vspace{-5pt}$ (10)
@@
Correspondingly, the other parts including both skill abstraction and low level execution are task-agnostic. For the high-level skill abstraction module, we apply a vector quantization (VQ) loss to refine the skill predictor. This VQ loss, $\mathcal{L}_{\mathrm{VQ}}$, ensures the embeddings produced by the skill predictor closely match the skill set vectors, thereby improving the interpretability and consistency of the skill representations. Inspired by VQ-VAE [45], we formulate it as:
$\mathcal{L}_{\mathrm{VQ}}=\mathbb{E}_{\boldsymbol{\tau}}\left[\|\mathbf{q}(\tilde{\boldsymbol{z}})-\tilde{\boldsymbol{z}}\|_{2}^{2}\right],\vspace{-5pt}$ (11)
@@
Additionally, we utilize a pre-trained distilBERT [39] as our language encoder, adopting the configuration consistent with LOReL [29] and LISA [13], while freezing its parameters to guarantee stability in language understanding. And we employ diverse settings to serve as the visual encoder to ensure fair comparison in different datasets. We elaborate the details in corresponding parts of Sec. 5. More training details are shown in Appendix G and we also provide some pseudo-code of our algorithm in Appendix B.
@@
Table 1: Task-wise success rates on LOReL Sawyer Dataset. We show our success rates compared to random policy, language-conditioned imitation learning (LCBC), language-conditioned Q-learning (LCRL), a flat non-hierarchical Decision Transformer (Lang-DT), and LISA. The results on each dataset are calculated over 3 seeds. SkillDiffuser outperforms all other methods in terms of average performance over all tasks. Best methods and those within 6% of the best are highlighted in bold.
@@ 5.1 Datasets
LOReL Sawyer Dataset [29] which is abbreviated from Language-conditioned Offline Reward learning, is composed of pseudo-expert trajectories or play data gathered from an arbitrary reinforcement learning policy, annotated with post-hoc crowd-sourced language directives. The LOReL Sawyer dataset encompasses 50k trajectories, each with 20 steps, in a simulated Sawyer robot environment. We assess our approach using the same six tasks as the original paper [29] with paraphrased instructions under five different situations (i.e. seen, unseen verb, unseen none, unseen verb + noun and human provided). This comes to a total of 77 instructions for all 6 tasks combined. More details about the dataset can be found in Appendix D.1.
@@ 5.2 Evaluation Results on LOReL Sawyer Dataset
Baselines. We employ random policy, language conditioned imitation learning (LCBC) [42], language conditioned Q-learning (LCRL) [20], Lang-DT (also known as Flat Baseline in [13]) and state-of-the-art skill-learning method LISA [13] as our baselines. We follow the same settings as LOReL [29] for the first three algorithms, and follow LISA [13] for the last two. The random policy serves as a baseline. And LCBC mimics offline dataset behaviors based on instructions, aligning with previous works focusing on imitation learning to achieve language conditioned behavior. In contrast, LCRL employs reinforcement learning, labeling each episode’s final state with language-instructed rewards. Lang-DT plays as a non-hierarchical benchmark with a Causal Transformer [25], while LISA incorporates a dual-transformer structure, with one for skill prediction and the other for action planning. We do not compare with LOReL planner [29] as it uses MPC on a learned reward function relying on human annotations, while LISA and ours learn with trajectory data only.
@@
Method Lang DT LOReL [29] LISA [13] SkillDiffuser
Success Rate 13.33 ±$\pm 1.3$ 18.18 ±$\pm 1.8$ 20.89 ±$\pm 0.6$ 25.21 ±$\pm 2.7$
Table 3: Performance on LOReL multi-step composition tasks.
5.3 Performance on LOReL Compositional Tasks
Settings. We conduct experiments following the same settings of unseen composition tasks of LISA [13] with 12 composition instructions in Tab. 3. Detailed instructions are listed in Appendix D.2 with such an example that “open drawer and move black mug right”. We extend the max number of episode steps from customary 20 to 40, as LISA.
Results. We observe SkillDiffuser achieves 2x the performance of non-hierarchical baseline (i.e. w/o skill abstraction) and also improves about 25% over LISA, highlighting its effectiveness. MPC-based LOReL planners are unable to perform as well in open scenarios like composition tasks.
5.4 Ablation Study on Meta-World Dataset
Settings. We conduct experiments on Meta-World MT10 benchmark with finely annotated instructions. We also use visual observations only. Details are in Appendix D.3.
Baselines. We evaluate our method against three baselines, all modified from existing models. The first, Flat R3M, is adapted from R3M [30] paper’s planner. As the original one utilizes the first four terms of robot proprioception, we eliminate them and make the planner focus exclusively on visual observations. The second baseline is a variant of our SkillDiffuser, lacking the high-level skill abstraction module but retaining the low-level conditional diffusion-based planner, to assess the impact of skill abstraction. This version integrates a two-layer MLP to predict options from visual and language inputs, functioning as a language-conditioned diffusion planner. The last baseline is our re-implemented LISA [13] for Meta-World Dataset to validate the efficiency of diffusion model. To ensure fairness, we use R3M as the visual encoder for all of these methods on Meta-World.
Method Lang Skill Set Diffuser Performance
@@ 5.4 Ablation Study on Meta-World Dataset
Table 4: Ablation study of language skill conditioning on Meta-World. Results of Flat R3M, Language-condition diffuser, LISA and our SkillDiffuser are shown here. All are averaged over 3 runs.
@@
Ablation on Language Skill Conditioning. Table 4 indicates our Meta-World MT10 task is quite different from and much more challenging than previous tasks that use states as observations [49] or take into robot proprioception [30]. We only utilize single-frame visual input and instructions. The Flat R3M method, lacking language conditioning and skill sets can only succeed on tasks like drawer-close and reach through behavior cloning. Lang-conditioned Diffuser and modified LISA both outperform Flat R3M, suggesting the value of each corresponding module. Our SkillDiffuser, discretizing skills into a skill set, achieves a 6% higher performance than language-conditioned diffuser and a 9.5% higher than LISA, demonstrating the effectiveness of this combinational architecture.
5.5 Ablation Study on Reusability of Learned Skills
To evaluate the reusability of our learned skills, we calculate the average number of different skills used for a single instruction and the total number of instructions using each skill of LOReL Sawyer Dataset in Table 5. (With max episode step being 20, we experiment with skill horizon 10.) We observe each single instruction uses 1.55 sub-skills on average and each skill is called multiple times than the number of instructions (75 with 5 eval episodes), verifying the transferability of learnt skills. As suggested in Tab.10 of [13], except a very small horizon will hurt the performance, learning sub-skills to get refined semantics helps perform different actions at different stages. Besides, we also visualize resulting images from applying discrete skills in Appendix E.1 to further validate skills’ interpretability.
5.6 Visualization Results of Learned Skill Set
@@
We show the visualization of skill set on LOReL compositional tasks here in Fig. 4 and results on original LOReL dataset in Fig. 5 and Meta-World dataset in Fig. 6 in Appendix C. The visual analysis of our SkillDiffuser’s learned skills on LOReL compositional tasks reveals that out of the 20-size skill set, our method learned 11 skills (e.g. pull handle [skill 0], open counter [skill 14], etc.) notably distinguished by their unique word highlights. These bright spots across eleven columns (changing from only one column at initial which corresponds to default BC) in the heatmap underscore the model’s ability to identify and isolate distinct skills from visual inputs, without an explicitly defined skill library. This indicates not only a significant interpretative advancement over previous diffusion-based planning but a successful abstraction of high-level skill representations.
# of learnt skills # of inst # of success Use 1 skill Use 2 skills Average
17 375 144 64 80 1.55
Freq of 17 skills 30, 8, 14, 10, 5, 19, 4, 7, 20, 2, 9, 20, 25, 6, 10, 3, 18
Table 5: Average number of different skills used for a single instruction and total number of instructions used for each skill.
@@
Figure 4: Visualization of skill heat map on LOReL Sawyer compositional tasks. We display the word frequency associated with a skill set of size 20 in LOReL, normalized by column. The data’s sparsity and distinct highlights indicate certain language tokens are uniquely linked to specific skills. There are eleven skills learned by our method. (zoom in for best view)
@@ 6 Conclusion
This paper presents SkillDiffuser, an integrated framework that enables robots to perform tasks from natural language instructions by enabling interpretable skill learning and conditional diffusion planning. It employs vector quantization to learn discrete and comprehensible skill representations directly from visual and linguistic demonstrations. Subsequently, these skills condition a diffusion model to generate state trajectories adhering to the learned skills. Through integrating hierarchical skill decomposition with conditional trajectory generation, SkillDiffuser can comprehend and execute abstract instructions for various manipulation tasks. Extensive experiments on manipulation benchmarks demonstrate state-of-the-art performance, highlighting its effectiveness for multi-step composition tasks and ability to automatically learn interpretable skills.
Acknowledgements
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622.
@@ Appendix B Pseudo-code of Training SkillDiffuser
Figure 6: Visualization of skill heat map on Meta-World Multi-Task 10 (MT10). There are eight skills learned by our method. (zoom in for best view)
@@ Appendix C More Visualizations
C.1 Visualization Results of Learned Skill Set
As mentioned before, we show the visualization results of skill set on LOReL Sawyer Dataset in Fig. 5 and Meta-World Multi-Task 10 (MT10) in Fig. 6. The visualization results show that out of a 20-size skill-set, our SkillDiffuser learned 11 skills for LOReL (e.g. pull drawer handle [skill 0], shut close container drawer [skill 15], etc.) and 8 skills for Meta-World MT10 (e.g. open push window [skill 0], open door with revolving joint [skill 16], etc.). The results demonstrate strong skill abstraction abilities. For example, the skill “shut close container drawer” abstracts different expressions like “shut drawer”, “shut container” into one skill semantic. In the heatmap, the presence of distinct bright spots across eleven columns strongly reaffirms the model’s capability to discern and pinpoint specific skills from visual inputs, in the absence of a pre-defined skill library. This observation is not just a testament to the model’s enhanced interpretative prowess over conventional diffusion-based planning approaches but also marks a remarkable stride in abstracting high-level skills into representations that are intuitively understandable by humans. Such evidence further validates the model’s proficiency in sophisticated skill identification and representation.
@@ D.1 LOReL Sawyer Dataset
Task Description
Closing the Drawer Involves the robot’s precise manipulation of a drawer to close it, testing spatial dynamics understanding and fine motor control.
Turning the Faucet Left Assesses the robot’s precision in rotational movements for turning a faucet to the left, a nuanced everyday action.
Pushing the Black Mug Right Requires the robot to push a specific object (black mug) to the right, testing its skills in object recognition and directional movement.
Table 6: Overview of tasks in LOReL Sawyer Dataset.
@@
D.2 LOReL Composition Tasks
We follow the same settings as LISA [13] to create 12 new composition tasks through combining original evaluation instructions as shown in Tab. 7.
Instructions
open drawer and move black mug right
pull the handle and move black mug down
move white mug right
move black mug down
close drawer and turn faucet right
close drawer and turn faucet left
turn faucet left and move white mug down
turn faucet right and close drawer
move white mug down and turn faucet left
close the drawer, turn the faucet left and move black mug right
open drawer and turn faucet counterclockwise
slide the drawer closed and then shift white mug down
Table 7: LOReL composition tasks
Additionally, we also incorporate tasks such as “move white mug right” and “move black mug down” to explore the composition of skills related to colors (e.g., black and white) and directions (e.g., right and down). This aims to explore whether such skills can be combined to fulfill complex instructions.
@@ D.2 LOReL Composition Tasks
Table 8: Annotated instructions for Meta-World MT10 tasks.
D.3 Meta-World Dataset
We sample 100 trajectories for each task of MT10 and form the expert dataset of 1000 trajectories. We have released our dataset with image observations on https://skilldiffuser.github.io.
Figure 10: Resulting images from applying skill 11 of Fig. 5. The black dashed line is a horizontal reference and please pay attention to the red oval region. (zoom in for best view)
Appendix E More Ablations
E.1 Ablation Study on Skill Interpretability
Resulting Images from Applying Discrete Skills We visualize resulting images from applying skill 11 of Fig. 5 which has grounding of “open, drawer, pull, dresser, container” (ranked from high frequency to low ones), consistent with its actual actions in Fig. 10. We can clearly observe a behavior of pulling the drawer. And we would like to clarify not all skills have clear semantic or action correspondences, while some do.
E.2 Ablation Study on Condition Guidance Weight
Classifier-free guidance is widely used in generative model domain for its ability to act as temperature control when setting guidance weight above 1 during inference. In all of our experiments, we set guidance weight to 1.2 by default. But we also conduct ablation study on the condition guidance weight here in Tab. 9. From the results, we find the guidance weight slightly greater than 1 helps the planner’s performance, while excessive weight hurts.
Guidance Weight 1.0 1.2 1.8 3.0 5.0
Success Rate on Seen Tasks 39.33% 46.67% 38.86% 39.03% 33.50%
Table 9: Ablation on guidance weight. (5 episodes over 3 seeds.)
Figure 11: Task-wise success rates (in %) on LOReL Sawyer Dataset.
Figure 14: Task-wise success rates (in %) on Meta-World MT10 Dataset.
Appendix F More Results
F.1 Task-wise Performance on LOReL Dataset
We further demonstrate the performance of our method and other baselines on LOReL Sawyer dataset in Fig 11 and 12. As can be seen from the figures, especially from Fig. 12, our method’s average performance on 5 rephrases is nearly 10 percentage points higher than the previous SOTA, which demonstrates its strong robustness against ambiguous language instructions.
F.2 Task-wise Performance on Meta-World
Appendix G Implementation Details
G.1 Hyper-parameters
@@ G.2 Architecture Details
G.3 Training Details